OpenStudy (anonymous):

How would you find a zero for this equation? 5 + 2*t + 0.6*t*Power[E, -0.3*t^2]​

OpenStudy (anonymous):

\[5 + 2t + 0.6t * e^{-0.3t^{2}}​\]

OpenStudy (anonymous):

$$5+2t+0.6t e^{-0.3t^2}=0$$ for sufficiently large \(|t|\) note that \(0.6 te^{-0.3t^2}\approx 0\) so we expect a solution close to that of \(5+2t=0\implies t=-2.5\) -- WolframAlpha confirms such a root \(t\approx -2.50144\)

OpenStudy (anonymous):

we know this because \(0.6(-2.5) e^{-0.3(-2.5)^2}=-1.5e^{-1.875}\approx -0.2300\)

OpenStudy (anonymous):

I was just wondering to myself if -2.50144 qualified as sufficiently large t, lol

OpenStudy (anonymous):

in practice, this does not have an exact analytic solution in terms of elementary functions; instead, you could use a fix-point algorithm like Newton's method to iteratively narrow in on a zero

OpenStudy (anonymous):

regardless, that's close enough and using the approximate function gives a good initial estimate for using an iterative algorithm like Newton's method

OpenStudy (anonymous):

cool, thanks

OpenStudy (anonymous):

much appreciated..

OpenStudy (anonymous):

http://ideone.com/ZUTfnT here's an example with dichotomic search

OpenStudy (anonymous):

That's pretty much the method I used.. a binary search doing a test for the sign at the midpoints.